What mass will push on a lever 80cm long. If the moment of 56.96Nm is caused about it's centre shaft?

1
Expert's answer
2021-12-23T10:48:28-0500

"\tau =F\cdot d=mg\cdot d\to m=\tau/(g\cdot d)=56.96/(9.8\cdot 0.8)="


"=7.3\ (kg)" . Answer